OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A. w + 5 = 116 B. 4w + 10 = 116 C. 2w + 5 = 116 D. 2w + 10 = 116

OpenStudy (anonymous):

B 4w+10=116 4w=116-10 w=106/4 w=26.5 Now check 26.5*2 +(26.5+5)*2 =53+31.5*2 =53+63 =116